One side of a rectangle is 4 feet longer than the other side. If the rectangle has an area of 96 square feet, find its dimensions
Answer by checkley75(3666)

x(x+4)=96
x^2+4x-96=0
(x-8)(x+12)=0
x-8=0
x=8 ans for one side.
8+4=12 for the other side.
Proof:
8*12=96
96=96